The gas tank of a truck is a cylinder 3.5 ft long with a diameter of 1.5 ft. At the gas station, a pump pours gas at the rate of 5 ft^3 per minute. How many minutes does it take to fill the empty tank with that pump? use the value 3.4
Answer by TimothyLamb(4379)

v = pi * r^2 * h
---
v = 3.14 * 0.75^2 * 3.5 = 6.19 cu.ft
---
r = v / t
t = v / r
---
r = 5 cu.ft/min
t = 6.19 cu.ft / 5 cu.ft/min = 1.24 minutes = 1 minute and 14.2 seconds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php